rapid walking through a distance of 2.2 km for 30 minutes per day requires 41 J while swimming for the same period of time through the same distance requires 226 J. Will these forms of exercise allow you to lose weight equivalent to 1 g of glucose (MW= 180 g/mol) in 1 week? the standard enthalpy of combustion is -2806 kJ.


1
Expert's answer
2022-01-13T06:19:01-0500

Moles of glucose = 1 / 180 = 0.00555 mole.


Total energy for 1 g is 2806 × 0.0055 = 16 kj approx .


So , yes these form of exercise allow you to lose weight .